find angle A if it supplements angle B which is equal to 70-x.
Answer by nerdybill(7384)

find angle A if it supplements angle B which is equal to 70-x.
.
First, you must realize that if two angles are supplementary, the sum of the two angles equals 180 degrees.
.
Angle B + Angle A = 180
Angle A = 180 - Angle B
Angle A = 180 - (70-x)
Angle A = 180 - 70 + x
Angle A = 110 + x (this is what they're looking for)